OpenStudy (anonymous):

A child is running on a moving sidewalk in an airport. When she runs against the sidewalk's motion, she travels 39 ft in 13 seconds. When she runs with the sidewalk's motion, she travels 65 ft in the same amount of time. What is the rate of the child on a still sidewalk and what is the rate of the moving sidewalk?

OpenStudy (anonymous):

is this for an algebra class?

OpenStudy (anonymous):

if so i think you just need to set up a system of equations. if its more complicated than that then im not sure

OpenStudy (anonymous):

let x be child speed and s be sidewalk speed

OpenStudy (anonymous):

distance rate and time

OpenStudy (anonymous):

x - s = 39

OpenStudy (anonymous):

x + s = ?

OpenStudy (anonymous):

solve for x and s?

OpenStudy (anonymous):

65

OpenStudy (anonymous):

i honestly dont know if thats right, but thats my guess

OpenStudy (anonymous):

by "sidewalk", do you mean a conveyor belt? A sidewalk is a motionless piece of cement

OpenStudy (anonymous):

yes 65 is correct

OpenStudy (anonymous):

do you know how to solve a system?

OpenStudy (anonymous):

belt

OpenStudy (anonymous):

2 equations 2 unknowns

OpenStudy (anonymous):

@billj5 you are doing absolutely right, we just did a question like that but Faine is having a little trouble so she's doing more, I hope you can help her though

OpenStudy (anonymous):

@Nishant_Garg ah good, ty b/c i wasnt sure @Faine so what are the two equations we have now?

OpenStudy (anonymous):

actually \[x-y=3\]and\[x+y=5\]

OpenStudy (anonymous):

all speeds are in ft/second

OpenStudy (anonymous):

ah ok, you divided by 13,

OpenStudy (anonymous):

apologize

OpenStudy (anonymous):

yeah, its an equation of speeds, so we would need to calculate the speeds first 39/13 and 65/13

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

@Faine what method do you prefer to use to solve this system?

OpenStudy (anonymous):

s - f = 3 s + f = 5

OpenStudy (anonymous):

5 = 3 = 8 2/8 = 4

OpenStudy (anonymous):